OpenStudy (anonymous):

write the decimal as a fraction or a mixed number in simplest form. 3.4

OpenStudy (anonymous):

34/10=17/5

OpenStudy (anonymous):

17/5=3 2/5

OpenStudy (anonymous):



OpenStudy (anonymous):

your 're right answer. thank you

OpenStudy (phi):

or you could say 3.4 is 3+ 0.4 so the answer will be 3 plus a fraction the fraction is 4/10 which simplifies (divide top and bottom by 2) to 2/5 so, as a mixed number, the answer is \[ 3 \frac{2}{5} \]